A surface area measuring 0.0024m^2 is parallel to an electric field with a magnitude of 5×〖10〗^4 N/C. Calculate the flux of the surface.
1
Expert's answer
2020-03-11T14:26:24-0400

"\Phi=B\cdot S\cdot \cos\alpha" .


According to the condition of the problem, the surface is parallel to the electric field lines. So, "\alpha=90°". We have


"\Phi=5\cdot10^4\cdot 0.0024\cdot\cos90°=0."